OpenStudy (zairahernandez):

How are the carbon, oxygen, and nitrogen cycles vital to sustaining life on Earth? a. Carbon, oxygen, and nitrogen are vital components of life on Earth. b. The carbon, oxygen, and nitrogen cycles allow vital elements to return to usable form by organisms. c. The carbon, oxygen, and nitrogen cycles are an important interface between biotic and abiotic factors. d. All of the above are reasons the carbon, oxygen, and nitrogen cycles are vital to life on Earth.

OpenStudy (abhisar):

Which do you think it is? Why or why not? It would be easier to tutor you if you explained what about the question is giving you trouble.

OpenStudy (zairahernandez):

I just simply dont get it lol I have tried google and everything it doesn't say anything in the book xc and im getting really upset.

OpenStudy (abhisar):

oh u don't need to get upset !

OpenStudy (abhisar):

\(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (zairahernandez):

Can It be d?

OpenStudy (abhisar):

Carbon cycle actually do the job of recycling.

OpenStudy (abhisar):

or any other biogeochemical cycle !

OpenStudy (anonymous):

I believe this is D, am I correct or no?

OpenStudy (anonymous):

Please correct me if I am wrong.

OpenStudy (zairahernandez):

Im saying d to LOL

OpenStudy (anonymous):

I am saying to be logically correct.

OpenStudy (abhisar):

They recycle elements like carbon, oxygen and nitrogen which are essential components of life. They link abiotic components like rock, soil etc with biotic components like animals and plants.

OpenStudy (abhisar):

So all the statement given is a valid reason. So yes D is the correct option !

OpenStudy (anonymous):

Thank you for clearing that up.

OpenStudy (zairahernandez):

Thank you guys (: @Abhisar can you please give a medal to Penguin?(:

OpenStudy (abhisar):

Here u go !

OpenStudy (anonymous):

Thank you.

OpenStudy (abhisar):

Lol ! @PenguinRulez give it to @ZairaHernandez

OpenStudy (zairahernandez):

thank you guys (:

OpenStudy (anonymous):

Your welcome